# VideoDirector: Precise Video Editing via Text-to-Video Models

### v2

@@ 1 Introduction
In this paper, we attribute these failures to: 1) Tightly spatial-temporal coupling. The entanglement of temporal and spatial (appearance) information in T2V models prevents vanilla pivotal inversion from compensating for the biases introduced by DDIM inversion. 2) Complicated spatial-temporal layout. The vanilla cross-attention control is insufficient to maintain the complex spatial-temporal layout of video content, resulting in low-fidelity editing results. By revisiting the fundamental mechanisms of the editing paradigm in T2V models, we argue that vanilla classifier-free guidance and null-text embeddings struggle to distinguish between temporal and spatial cues. Consequently, they fail to compensate for the biases introduced by DDIM inversion, resulting in meaningless latents. In addition, the temporal layers in T2V models build a complicated relationship between the spatial-temporal tokens. As a result, the latents are vulnerable to the crosstalk introduced by cross-attention manipulation.
@@
  • We develop a self-attention control strategy to maintain the complex spatial-temporal layout and enhance fidelity.
@@ 2 Related Work
Text-to-Image Editing Recent advances in T2I generation models have promoted the rapid development of text-guided image editing methods [23, 8, 20, 19, 24, 26, 29]. Hertz et al. [8] introduced Prompt-to-Prompt to edit images via DDIM inversion and manipulation of cross-attention maps. Specifically, techniques such as Word Swap, Phrase Addition, and Attention Re-weighting are performed to modify the attention maps based on text prompts. Since the DDIM inversion introduces biases by approximating noise latent***For more information on DDIM inversion, please refer to our supplementary material., Mokady et al. [20] introduced a step-wise null-text embedding $\phi_{t}$ optimized after DDIM inversion for compensation. This optimization refines the denoising trajectory by compensating for DDIM inversion biases, enhancing both reconstruction quality and editing precision. Different from this pipeline, DreamBooth [24] fine-tuned a pre-trained T2I model [25] to synthesize subjects in prompt-guided diverse scenes using reference images as additional conditions.
Text-to-Video Editing Numerous efforts have been made to extend T2I models directly to video editing [16, 3, 5, 12]. Tune-A-Video [28] developed a tailored spatio-temporal attention mechanism and an efficient one-shot tuning strategy to tune an input video. Video-P2P [16] transforms a T2I model to a video-customized Text-to-Set (T2S) model through fine-tuning to achieve semantic consistency across adjacent frames. TokenFlow [5] explicitly propagates token features based on inter-frame correspondences using the T2I model without any additional training or fine-tuning. RAVE [12] utilizes Controlnet and introduces random shuffling of latent grids to ensure temporal consistency. Flatten [3] incorporates optical flow into the attention module of the T2I model to address inconsistency issues in text-to-video editing. Due to the lack of temporal generation capacity of T2I models, the aforementioned methods still suffer from results with inferior temporal coherence, realism, and motion smoothness.
@@ 3 Method
3.1 Problem Definition & Challenge Discussion
Given an input video $V_{i}\in\mathbb{R}^{H\times W\times F}$, a descriptive prompt $\mathcal{C}$ (“A wolf turns its head, with many trees in the background”), and an editing prompt $\mathcal{C}^{e}$ (replacing “wolf” with “husky”), the objective of video editing is to obtain an edited target video $V_{o}$ using a generation model $G$:
$V_{o}=\mathcal{D}\left(V_{i}~{}|~{}G,(\mathcal{C},\mathcal{C}^{e}),\mathcal{R}\right).$ (1)
Here, $G$ refers to a T2I or T2V model and $\mathcal{R}$ denotes an optional regularization term obtained from external models. Intuitively, the edited videos should be of high quality in terms of the following four aspects: (1) Accuracy: The wolf is accurately replaced by a husky, which can be evaluated using Pick score [13]. (2) Fidelity: The backgrounds are well preserved, which can be measured by masked PSNR and LPIPS. (3) Motion Smoothness: The husky mimics the motion of the wolf with high smoothness, which can be assessed using VBench [11]. (4) Realism: The husky is enriched with realistic, hallucinated details consistent with real-world physical laws, such as its breathing, leaves swaying in the wind, and sunlight filtering through the leaves.
Currently, most video editing methods employ T2I models as $G$ and rely on external regularizations (e.g., optical flow, depth maps) as $\mathcal{R}$ to incorporate temporal information. However, since T2I models suffer limited temporal generation capacity and the additional regularization delivers insufficient temporal cues of the edited contents, these methods fall short in motion smoothness and realism.
In this paper, we argue that incorporating T2V models is the key to address the above issues. However, directly extending the typical “inversion-then-editing” paradigm to T2V models faces critical challenges. First, the vanilla diffusion pivotal inversion [20] fails to accurately reconstruct the input video. Second, prompt-to-prompt [8] editing cannot well preserve the unedited contents. To remedy this, we propose a spatial-temporal decoupled guidance module and multi-frame null-text optimization to accomplish pivotal inversion for the T2V model, as detailed in Sec. 3.2. Additionally, we introduce a tailored attention control strategy to achieve precise editing while preserving the original, unedited content, as described in Sec. 3.3. Moreover, this mutual attention strategy enhances harmony, allowing the edited content to be seamlessly integrated, thereby improving the overall realism of the edited videos.
@@ 4.1 Evaluation
Quantitative Evaluation. We evaluate the edited videos based on four key aspects, as outlined in the editing quality objectives described in 3.1: Accuracy, Fidelity, Motion Smoothness, and Realism. For accuracy, we use the Pick score (PS) [13] to assess the alignment quality. For fidelity, we calculate the masked PSNR (m.P) and LPIPS (m.L) to evaluate the preservation quality of the original, unedited content. For motion smoothness (MS), we utilize VBench [11] to assess whether the motion in the edited video is smooth and adheres to real-world physical laws. We also conducted a user study (US) to evaluate the realism of the edited videos. Nine participants were asked to rank all competing methods from best (rank 1) to worst (rank 5) in terms of realism and editing effectiveness, and the mean score was calculated. As shown in Table 1, our method outperforms all other methods across all metrics, demonstrating superior quantitative editing performance.
@@ 4.2 Ablation study
Spatial-Temporal Decoupled Guidance. As shown in Fig. 9 and  Fig. 7, removing the STDG significantly degrades the performance of both reconstruction and video editing. This degradation is evident from the severe color flickering and unstable video quality observed. These findings highlight the critical role of the STDG in ensuring effective video reconstruction and editing.
@@ 5 Conclusion
[…]
Supplementary Material
I Preliminaries
Latent Diffusion Models (LDMs).
$\boldsymbol{z}_{t-1}=\sqrt{\alpha_{t-1}}\underbrace{\left(\dfrac{\boldsymbol{z}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\boldsymbol{z}_{t})}{\sqrt{\alpha_{t}}}\right)}_{\text{\normalsize{``predicted} }\boldsymbol{z}_{0}\text{\normalsize{''}}}+\underbrace{\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\boldsymbol{z}_{t})}_{\text{\normalsize{``direction pointing to} }\boldsymbol{z}_{t}\text{\normalsize{''}}}.$
$\mathcal{L}(\phi_{t})=\|z_{t-1}^{*}-z_{t-1}\|_{2}^{2},$ (13)
II Discussion about Null-text Optimization
Replacing the multi-frame strategy with a shared null-text embedding is effective for objects with minimal deformation, such as the “driving car" shown in Fig. I. In these cases, the STDG provides sufficient temporal and motion guidance. However, relying solely on the STDG leads to suboptimal reconstruction and editing results in videos with dynamic objects that undergo significant deformation, as illustrated in Fig. I. Multi-frame null-text optimization is crucial for videos featuring such dynamic objects. While the STDG offers global temporal and spatial guidance, the null-text embedding refines detailed motion and appearance information by building on the STDG and pivotal latent.
III Discussion about SAM2 Mask
In our method, we employ SAM2 [22] to distinguish the target objects for editing from the others. While the mask generated by SAM2 is able to segment fine structures, these rich details can make the editing process fragile and vulnerable to disruptions caused by segmentation masks, as shown in Fig. II. To mitigate this issue, we combine the SAM2 mask with an ellipse mask that is coarsely aligned with the SAM2 mask during the pivotal inversion and editing process. In this way, the combined mask enhances robustness of our method to mask disruptions and improves the harmony between the edited and the remaining contents, as illustrated in Fig. II.
IV Pseudo Code and More Results
The pseudo-code for our method is provided in Algorithm I. Descriptions of the variables used in the algorithm can be found in Sec. 3. Stage 1 corresponds to Sec. 3.2, and Stage 2 corresponds to Sec. 3.3. Here, $e_{t}^{*}$ denotes the DDIM sampling latents of the editing path in Stage 2.
More edited results are shown in Fig III to Fig IV, Fig V, and Fig VI. along with our editing prompts. Additionally, we provide an MP4 video in the supplementary material.
V Limitation
The edited videos in this paper are limited to 16 frames due to the high memory cost of the T2V model. In addition, we simultaneously sample two separate latent paths during editing. Our method consumes approximately 16GB more GPU memory usage compared to Video-p2p [16]. In the future, we will further focus on extending the method to handle longer video sequences.
Figure I: Shared Null-text optimization used for reconstruction and editing.
Figure II: Sam2 Mask combines the ellipse mask to enhance the editing robustness.
Figure III: More results.
Figure IV: More results.
Figure V: More results.
Figure VI: More results.
Algorithm I VideoDirector
1:Input: video ${V_{i}}\in\mathbb{R}^{{F}\times{H}\times{W}}$, regularization term $\mathcal{R}$: SAM2 masks ${\mathcal{M}}\in\mathbb{R}^{{F}\times{H}\times{W}}$ [22], original and editing prompts: $\mathcal{C}$ and $\mathcal{C}^{e}$, generation model G: T2V diffusion network $\epsilon_{\theta}$ [6].
2:Edited video ${V_{o}}\in\mathbb{R}^{{F}\times{H}\times{W}}$.
3:Stage 1: Video Pivotal Inversion
4:$z^{*}=\mathcal{E}(V_{i})$triangleright Encoder $\mathcal{E}(\cdot)$ convert the input video to latents.
5:for $t=0$ to $T$ do $\triangleright$ Iterate over $T$ timesteps.
6:       $\boldsymbol{z}_{t+1}^{*}=\sqrt{\alpha_{t+1}}\left(\dfrac{\boldsymbol{z}_{t}^{*}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\boldsymbol{z}_{t}^{*})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t+1}}\epsilon_{\theta}(\boldsymbol{z}_{t}^{*})$triangleright DDIM inversion.
7:end for
8:for $t=T$ to 00 do $\triangleright$ Iterate over $T$ timesteps in reverse.
9:       $\mathcal{T}_{+}=\epsilon_{\theta}^{(\mathcal{T})}(\boldsymbol{z}_{t}^{*},\mathcal{C},t),\quad\mathcal{T}_{-}=\epsilon_{\theta}^{(\mathcal{T})}(\boldsymbol{z}_{t},\mathcal{C},t)$triangleright Extract temporal features.
10:       $\mathcal{K}_{+}=\epsilon_{\theta}^{(\mathcal{K})}(\boldsymbol{z}_{t}^{*},\mathcal{C},t),\quad\mathcal{K}_{-}=\epsilon_{\theta}^{(\mathcal{K})}(\boldsymbol{z}_{t},\mathcal{C},t)$triangleright Extract spatial features.
11:       $\mathcal{L}_{\mathcal{T}}=\mathcal{M}^{f/b}_{\mathcal{T}}\cdot\mathcal{M}_{\mathcal{T}}\cdot\|{(\mathcal{T}_{+}-\mathcal{T}_{-})}\|_{2}^{2},\quad\mathcal{G}_{\mathcal{T}}^{f/b}=\dfrac{\partial({\mathcal{L}_{\mathcal{T}}})}{\partial z_{t}}$triangleright Temporal Guidance.
12:       $\mathcal{L}_{\mathcal{K}}=\mathcal{M}^{f/b}_{\mathcal{K}}\cdot\|{(\mathcal{K}_{+}-\mathcal{K}_{-})}\|_{2}^{2},\quad\mathcal{G}_{\mathcal{K}}^{f/b}=\dfrac{\partial({\mathcal{L}_{\mathcal{K}}})}{\partial z_{t}}$triangleright Spatial Guidance.
13:       $\mathcal{G}_{t}=\eta_{f}\cdot\mathcal{G}_{\mathcal{T}}^{f}+\eta_{b}\cdot\mathcal{G}_{\mathcal{T}}^{b}+\zeta_{f}\cdot\mathcal{G}_{\mathcal{K}}^{f}+\zeta_{b}\cdot\mathcal{G}_{\mathcal{K}}^{b}$triangleright Total Guidance.
14:       for $iter=0$ to $N$ do $\triangleright$ Iterative Null-text Optimize for $N$ steps.
15:             $\hat{\epsilon_{\theta}}=\epsilon_{\theta}(z_{t},\mathcal{C},t)+\omega[\epsilon_{\theta}(z_{t},\mathcal{C},t)-\epsilon_{\theta}(z_{t},\{\phi_{t}\},t)]$triangleright CFG.
16:             $\overline{\epsilon_{\theta}}=\hat{\epsilon_{\theta}}-(\sqrt{1-\alpha_{t}})\mathcal{G}_{t}$triangleright STDG, the guidance is applied following the formula (14) from [4].
17:             $\boldsymbol{z}_{t-1}=\sqrt{\alpha_{t-1}}\left(\dfrac{\boldsymbol{z}_{t}-\sqrt{1-\alpha_{t}}\hat{\epsilon_{\theta}}}{\sqrt{\alpha_{t}}}\right)+(\sqrt{1-\alpha_{t-1}})\overline{\epsilon_{\theta}}$triangleright DDIM sampling.
18:             ($\mathcal{L}(\{\phi_{t}\})=\|z_{t-1}^{*}-z_{t-1}\|_{2}^{2}$triangleright Null-text Optimize.
19:       end for
20:end for
21:Stage 2: Attention Control for Video Editing
22:for $t=T$ to 00 do $\triangleright$ DDIM sampling.
23:       for $l=0$ to $L$ do $\triangleright$ Pass through the U-Net of the T2V model.
24:             $Q_{t}^{(l)}=\epsilon_{\theta}^{(l)(Q)}(z_{t}^{*}),\quad K_{t}^{(l)}=\epsilon_{\theta}^{(l)(K)}(z_{t}^{*}),\quad V_{t}^{(l)}=\epsilon_{\theta}^{(l)(V)}(z_{t}^{*})$triangleright Extract $Q$, $K$, $V$ of reconstruction path.
25:             $Q_{t}^{*(l)}=\epsilon_{\theta}^{(l)(Q)}(e_{t}^{*}),\quad K_{t}^{*(l)}=\epsilon_{\theta}^{(l)(K)}(e_{t}^{*}),\quad V_{t}^{*(l)}=\epsilon_{\theta}^{(l)(V)}(e_{t}^{*})$triangleright Extract $Q$, $K$, $V$ of editing path.
26:             if $SelfAttention$ then $\triangleright$ Self Attention Control.
27:                    $\widehat{Attn}=\begin{cases}W_{t}^{(l)}\cdot V_{t}^{*(l)},\quad\text{if }t<\tau_{s},\vspace{0.2cm}\\ S\left(\displaystyle\frac{Q_{t}^{*(l)}\cdot\hat{K}_{t}^{\top}}{\sqrt{d}}\otimes\left[\mathbf{1}\mid\mathcal{M}^{f}\right]\right)\cdot\hat{V}_{t},&\text{otherwise.}\\ \end{cases}$triangleright Calculate attention features in SA-I and SA-II.
28:             else if $CrossAttention$  then $\triangleright$ Cross Attention Control.
29:                    $M_{t}^{C(l)}=\begin{cases}\boldsymbol{C}\cdot[\boldsymbol{\gamma}\cdot(M_{t}^{*(l)})+(\boldsymbol{1}-\boldsymbol{\gamma})\cdot(M_{t}^{\prime(l)})],&\text{if}t<\tau_{c},\\ M_{t}^{*(l)},\quad\text{otherwise.}\end{cases}$triangleright Calculate Cross Attention Maps.
30:             end if
31:             Update edited latent $\epsilon_{\theta}^{(l)}(e_{t}^{*})$. $\triangleright$ This edited latent updating contains $\epsilon_{\theta}^{(l)}(e_{t}^{*},\mathcal{C},t)$ and $\epsilon_{\theta}^{(l)}(e_{t}^{*},\{\phi_{t}\},t)$.
32:       end for
33:       $\hat{\epsilon_{\theta}}=\epsilon_{\theta}(e_{t}^{*},\mathcal{C},t)+\omega[\epsilon_{\theta}(e_{t}^{*},\mathcal{C},t)-\epsilon_{\theta}(e_{t}^{*},\{\phi_{t}\},t)]$triangleright CFG.
34:       $\overline{\epsilon_{\theta}}=\hat{\epsilon_{\theta}}-(\sqrt{1-\alpha_{t}})\mathcal{G}_{t}$triangleright STDG.
35:       $e^{*}_{t-1}=\sqrt{\alpha_{t-1}}\left(\dfrac{e^{*}_{t}-\sqrt{1-\alpha_{t}}\hat{\epsilon_{\theta}}}{\sqrt{\alpha_{t}}}\right)+(\sqrt{1-\alpha_{t-1}})\overline{\epsilon_{\theta}}$triangleright DDIM sampling using edited latent.
36:end for
37:return $V_{o}=\mathcal{DE}(e_{0}^{*})$. $\triangleright$ Decoder $\mathcal{DE}(\cdot)$ convert edited latents into output video.

### v3

@@ 1 Introduction
In this paper, we attribute these failures to: 1) Tightly spatial-temporal coupling. The entanglement of temporal and spatial (appearance) information in T2V models prevents vanilla pivotal inversion from compensating for the biases introduced by DDIM inversion. 2) Complicated spatial-temporal layout. The vanilla cross-attention control is insufficient to maintain the complex spatial-temporal layout of video content, resulting in low-fidelity editing results.
@@
Our key insight is that accurate reconstruction is the basis of high-quality video editing, where accurate reconstruction (Sec. 3.2) refers to precisely resampling the input video from noise using T2V models, and video editing (Sec. 3.3) can be formulated as precisely controlling the reconstruction trajectory to achieve the desired modifications. Previously, directly integrating DDIM inversion-based methods with T2V models cannot produce satisfactory results Fig. 2(a). Our approach demonstrates the feasibility of controlling the denoising trajectory for precise editing via T2V model.
@@
  • We develop a self-attention control strategy to maintain the complex spatial-temporal layout and enhance fidelity. Our approach demonstrates the feasibility of controlling the denoising trajectory for precise editing.
@@ 2 Related Work
Text-to-Image Editing Recent advances in T2I generation models have promoted the rapid development of text-guided image editing methods [24, 9, 21, 20, 25, 27, 30]. Hertz et al. [9] introduced Prompt-to-Prompt to edit images via DDIM inversion and manipulation of cross-attention maps. Specifically, techniques such as Word Swap, Phrase Addition, and Attention Re-weighting are performed to modify the attention maps based on text prompts. Since the DDIM inversion introduces biases by approximating noise latent, Mokady et al. [21] introduced a step-wise null-text embedding $\phi_{t}$ optimized after DDIM inversion for compensation. This optimization refines the denoising trajectory by compensating for DDIM inversion biases, enhancing both reconstruction quality and editing precision. Different from this pipeline, DreamBooth [25] fine-tuned a pre-trained T2I model [26] to synthesize subjects in prompt-guided diverse scenes using reference images as additional conditions.
Text-to-Video Editing Numerous efforts have been made to extend T2I models directly to video editing [17, 5, 6, 13, 2, 4]. Tune-A-Video [29] developed a tailored spatio-temporal attention mechanism and an efficient one-shot tuning strategy to tune an input video. Video-P2P [17] transforms a T2I model to a video-customized Text-to-Set (T2S) model through fine-tuning to achieve semantic consistency across adjacent frames. TokenFlow [6] explicitly propagates token features based on inter-frame correspondences using the T2I model without any additional training or fine-tuning. RAVE [13] utilizes Controlnet and introduces random shuffling of latent grids to ensure temporal consistency. Flatten [5] incorporates optical flow into the attention module of the T2I model to address inconsistency issues in text-to-video editing. Due to the lack of temporal generation capacity of T2I models, the aforementioned methods still suffer from results with inferior temporal coherence, realism, and motion smoothness.
@@ 3 Method
3.1 Preliminaries
Latent Diffusion Models (LDMs).
$\boldsymbol{z}_{t-1}=\sqrt{\alpha_{t-1}}\left(\dfrac{\boldsymbol{z}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\boldsymbol{z}_{t})}{\sqrt{\alpha_{t}}}\right)+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}(\boldsymbol{z}_{t}).$
$\mathcal{L}(\phi_{t})=\|z_{t-1}^{*}-z_{t-1}\|_{2}^{2},$ (6)
@@ 4.1 Evaluation
Quantitative Evaluation. We evaluate the edited videos based on four key aspects: Accuracy, Fidelity, Motion Smoothness, and Realism. For accuracy, we use the Pick score (PS) [14] to assess the alignment quality. For fidelity, we calculate the masked PSNR (m.P) and LPIPS (m.L) to evaluate the preservation quality of the original, unedited content. For motion smoothness (MS), we utilize VBench [12] to assess whether the motion in the edited video is smooth and adheres to real-world physical laws. We also conducted a user study (US) to evaluate the realism of the edited videos. Nine participants were asked to rank all competing methods from best (rank 1) to worst (rank 5) in terms of realism and editing effectiveness, and the mean score was calculated. As shown in Table 1, our method outperforms all other methods across all metrics, demonstrating superior quantitative editing performance.
@@ 4.2 Ablation study
Abl. w/o STDG shared NT w/o CA w/o SA w/o SA-I w/o SA-II Ours
MS ↑ 89.23% 97.21% 96.58% 90.37% 93.50% 97.62% 97.68%
PS ↑ 20.39 20.44 21.06 20.10 20.67 20.63 21.64
m.P ↑ 19.09 19.01 21.13 14.87 16.93 20.27 21.37
m.L ↓ 0.369 0.346 0.301 0.537 0.418 0.371 0.270
Table 2: Quantitative comparison of modules in video editing.
@@
Spatial-Temporal Decoupled Guidance. As shown in Fig. 9 and  Fig. 7, removing the STDG significantly degrades the performance of both reconstruction and video editing. This degradation is evident from the severe color flickering and unstable video quality observed. These findings highlight the critical role of the STDG in ensuring effective video reconstruction and editing. The quantitative results of the ablation study are shown in Tab. 2, demonstrating the effectiveness of our proposed modules.
@@ 5 Conclusion
6 Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (No. U20A20185, 62372491, 62301601), the Guangdong Basic and Applied Basic Research Foundation (No. 2022B1515020103, 2023B1515120087), the Shenzhen Science and Technology Program (No. RCYX20200714114641140), the Science and Technology Research Projects of the Education Office of Jilin Province (No. JJKH20251951KJ).